OpenStudy (anonymous):

WHAT IS (2x^-2)3 3x^-5

jhonyy9 (jhonyy9):

2x*-2 =1/2x*2 --- (2x*(-2))3=(1/2x*(2))3=3/2x*2 3x*(-5) =1/3x*5